(1 point) If k≥1, the graphs of y=sinx and y=ke−x intersect for x≥0. Find the smallest value of k for which the graphs are tangent.
1
Expert's answer
2020-03-20T10:45:27-0400

y=f'(x0) *(x-x0) +f(x0)

y=cos(x0) *(x-x0) +sin(x0)

cos(x0) =-1

x0=π+2πn

for n=0:

y=ke-π=sin(π) =0

k=π/e